OpenStudy (anonymous):

Help! I will give the best answer (as long as you show me how to do it) a medal and I will become a fan. 1) How many molecules of propane are required to produce 8.50 moles of water? 2) How may molecules of water are produced when 26.0 grams of oxygen are used? 3) How many molecules of water are produced from 12.0 moles propane? 4) How many molecules of oxygen are required to produce 5.00 grams of water?

OpenStudy (anonymous):

The given is C3H8(g)+5O2(g)3CO2(g)+4H2O(g)

OpenStudy (abb0t):

All of these are what you would call a \(\sf \color{Red}{combustion ~reaction}\), meaning, you're taking some organic molecule in the presence of oxygen to give carbon dioxide and water.

OpenStudy (anonymous):

The box is an arrow to the right

OpenStudy (anonymous):

It is stoichiomentry

OpenStudy (abb0t):

Now, if you're given 8.50 molecules, that means you need to convert from moles of H\(_2\)O \(\rightarrow\) molecules of propane. Yes, this is stoichiometry. So you would follow this procedure: \(\sf \color{blue}{8.50~mol~H_2O \times \frac{1~mol~propane}{4~mol~H_2O} = mol~of~propane}\) now, you need to apply avogadro's constant, which is 6.022 \(\times\) 10\(^{23}\) \(\sf \frac{molecules}{mol}\)

OpenStudy (abb0t):

You would follow the same steps for the others. This is called \(dimensional~analysis\).